Circular No. 6383

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1739–278 P. Durouchoux, Service d’Astrophysique, Saclay; I. A. Smith, Rice

University; K. Hurley, University of California, Berkeley; A. S. B. Schultz, NASA Ames Research Center; L. B. F. M. Waters, University of Amsterdam (UA); J. van Paradijs, University of Alabama at Huntsville and UA; and P. Wallyn, Jet Propulsion Laboratory, report: “We observed the GRS 1739–278 region (IAUC 6348) with the SEST 1.3-mm bolometer (center frequency 235 GHz; band 50 GHz; beam 23′′) on Mar. 25.42 UT. The antenna was pointed toward the source direction given by Borozdin it et al. (IAUC 6350). The pointing procedure was as follows: 10 s in the source direction, followed by 10 s off-source (70′′ offset in azimuth in one direction) and 10 s off-source (70′′ offset in the opposite azimuthal direction). We accumulated a total of 2400 s on-source, and used Uranus for calibration. The atmospheric transmission was very good (τ = 0.06, corresponding to an atmospheric conversion factor of ∼ 7 percent). We detected a possible source, with a flux density of 17.6 ± 5.3 mJy. Dividing the 2400 s on-source observation into three sets of 800 s suggests that the source is variable: Mar. 25.445–25.476, 31.0 ± 8.9 mJy; Mar. 25.476–25.493, 17.0 ± 8.8 mJy; Mar. 25.494–25.526, 1.25 ± 10.2 mJy. This variability suggests that we located the source, although we caution that our beam did not cover the entire x-ray error box. We encourage further multiwavelength observations to confirm and monitor this potential counterpart.”

R. M. Hjellming and M. P. Rupen, National Radio Astronomy Observatory; and J. Marti, University of Barcelona; F. Mirabel, Saclay; and L. F. Rodriguez, Universidad Nacional Autonoma de México, report the VLA detection of a probable radio counterpart to GRS 1739–278, discovered by the SIGMA/Granat team (IAUC 6348): “With a position of α = 17^{h}42^{m}40^{s}.03 ± 0^{s}.02, δ = −27^{o}44′52′′.7 ± 0′′.3 (equinox 2000.0), the radio source is just inside the 1′ error radius of the MIR-KVANT-TTM team’s (IAUC 6350) position. Unlike four other radio sources in the 4.9-GHz field of observation, which are outside of the 1′ error region and which appear stable in radio flux, this radio source was 1.1, 1.5, 3.7, and 4.7 mJy at 4.9 GHz on Mar. 24, Apr. 8, 14, and 18, respectively, indicating a timescale of variations typical of the synchrotron radiation sources associated with x-ray transients and x-ray binaries.”

1996 April 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green